OpenStudy (albert0898):

Ticket sales for a music concert totaled $2,160. Three times as many tickets were sold for the Saturday night concert as were sold for the Sunday afternoon concert. Two times as many tickets were sold for the Friday night concert were sold for the Saturday afternoon concert. Tickets for all three concerts sold for $2.00 each. Find the number of tickets sold for the Saturday night concert. *ATTN: Please explain all the steps, thanks*

OpenStudy (zepp):

Do you know how to start this off?

OpenStudy (albert0898):

Well I know how to start off similar problems. It's just I don't get how to set up an equation from this information... "Three times as many tickets were sold for the Saturday night concert as were sold for the Sunday afternoon concert. Two times as many tickets were sold for the Friday night concert were sold for the Saturday afternoon concert."

OpenStudy (zepp):

Ok we have two days here Let number of tickets sold Saturday be x and Sunday be y

OpenStudy (zepp):

They say: 2$ per ticket for both night.

OpenStudy (zepp):

3 times as many tickets were sold for the Saturday night concert as were solve for the Sunday afternoon. That mean that y=3x

OpenStudy (anonymous):

first you need to get the total number of tickets sold which is 2160/2=1080 tickets

OpenStudy (zepp):

Therefore we can build up an equation ( Tickets sold Saturday + Tickets sold Sunday ) * $2 = $2,160

OpenStudy (zepp):

Tickets sold Saturday = x Tickets sold Sunday = 3x (x+3x) * $2 = $2,160 4x * $2 = $2,160

OpenStudy (zepp):

Can you take it from there?

OpenStudy (albert0898):

Here's what my book says... $2 * x + $2 * 3x +2 * 2x = $2160 2x + 6x + 4x = $2160 12x = 2160 Divide by 12 x = 180 Can you please explain that to me ...

OpenStudy (zepp):

Darn, I forgot there's Friday's concert, too :x

OpenStudy (albert0898):

:c

OpenStudy (zepp):

But we can correct our mistake!

OpenStudy (zepp):

Two times as many tickets were sold for the Friday night concert were sold for the Saturday afternoon concert. There, they say and Friday night = Saturday's concert * 2

OpenStudy (zepp):

Therefore, Tickets sold Friday = 2x Tickets sold Saturday = x Tickets sold Sunday = 3x Let's take the equation we build up earlier; ( Tickets sold Friday + Tickets sold Saturday + Tickets sold Sunday ) * $2 = $2,160 (2x + x + 3x) * $2 = $2,160 :)

OpenStudy (albert0898):

Ah, I see, thanks !

OpenStudy (albert0898):

Oh and the 'times 2' represents the 2 dollar fee per person, right?

OpenStudy (zepp):

Per ticket, yes.

OpenStudy (albert0898):

Thanks xD

OpenStudy (zepp):

yw :)